# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 2 Lagrangian multiforms on coadjoint orbits
In this section, we briefly review the results of [CDS] where a Lagrangian multiform was obtained which produces a hierarchy of equations of motion in Lax form as its multi-time Euler-Lagrange equations. We will, in fact, be interested in applying this framework in the infinite-dimensional setting for which we refer more precisely to [CDS, Section 7]. In order to avoid confusion with the finite-dimensional Lie algebra $\mathfrak{g}$ used throughout the rest of this paper, in this section we denote by $\mathbb{g}$ the infinite-dimensional Lie algebra of interest. The construction of the specific Lie algebra $\mathbb{g}$ relevant for the cyclotomic Gaudin model will be the subject of the next section.
The framework of Lie dialgebras for constructing hierarchies of compatible Lax equations involves a number of ingredients. For the purpose of this paper, we will actually only require a particular instance of this general framework corresponding to the so-called Adler-Kostant-Symes scheme [A, K, Sy], so we will mainly focus on this case to ease the comparison with later sections. The ingredients are:
@@
    $\langle\cdot,\cdot\rangle:V\times\mathbb{g}\longrightarrow\mathbb{C}$ (2.1)
    with some representation $V$ of $\mathbb{G}$, allowing one to use $V$ as a model for the dual space $\mathbb{g}^{\ast}$ of $\mathbb{g}$. We will denote the representation of $\mathbb{G}$ by $\operatorname{Ad}^{\ast}:\mathbb{G}\times V\to V$, $(\varphi,\xi)\mapsto\operatorname{Ad}^{\ast}_{\varphi}\xi$ and the corresponding Lie algebra representation by $\operatorname{ad}^{\ast}:\mathbb{g}\times V\to V$, $(X,\xi)\mapsto\operatorname{ad}^{\ast}_{X}\xi$, and refer to these as the coadjoint representations.
    Note that if $\mathbb{g}$ were finite-dimensional then a good choice of representation space $V$ would be the algebraic dual $\mathbb{g}^{\prime}$ of $\mathbb{g}$, in which case (2.1) is given by the canonical pairing. In the infinite-dimensional setting we are considering, the algebraic dual is too big. When $\mathbb{g}$ is equipped with a nondegenerate invariant symmetric bilinear form $(\cdot,\cdot):\mathbb{g}\times\mathbb{g}\to\mathbb{C}$ then one possible replacement for the algebraic dual is afforded by the smooth dual $(\mathbb{g},\cdot)\subset\mathbb{g}^{\prime}$ which is canonically isomorphic to $\mathbb{g}$ itself. However, in our present setting it turns out that the smooth dual will not be the appropriate notion of dual space, which is why we need the above more general working definition for the dual space $\mathbb{g}^{\ast}$.
@@
    A linear map $R:\mathbb{g}\to\mathbb{g}$ satisfying the modified classical Yang–Baxter equation (mCYBE)
    $[R(X),R(Y)]-R\left([R(X),Y]+[X,R(Y)]\right)=-[X,Y]\,,\quad\forall\leavevmode\nobreak\ X,Y\in\mathbb{g}\,.$ (2.2)
    This is the crucial ingredient providing us with the Lie dialgebra $(\mathbb{g},\mathbb{g}_{R})$ where $\mathbb{g}_{R}$ is the vector space $\mathbb{g}$ equipped with a second Lie bracket $[X,Y]_{R}=\frac{1}{2}\big{(}[R(X),Y]+[X,R(Y)]\big{)}$. In the Adler-Kostant-Symes scheme, this is determined by a direct sum decomposition
@@
    into complementary subalgebras $\mathbb{g}_{\pm}\subset\mathbb{g}$. A solution to (2.2) is then given by $R=P_{+}-P_{-}$ where $P_{\pm}:\mathbb{g}\to\mathbb{g}_{\pm}$ denote the projections onto these respective subalgebras. In this case, the Lie algebra $\mathbb{g}_{R}$ is the direct sum of Lie algebras $\mathbb{g}_{R}=\mathbb{g}_{+}\oplus\mathbb{g}_{-}$. Now (2.3) induces a corresponding direct sum decomposition of the model $V$ for the dual space $\mathbb{g}^{\ast}$, namely
    $V=V_{+}\dotplus V_{-}$ (2.4)
    where $V_{\pm}$ is the orthogonal complement of $\mathbb{g}_{\pm}$ with respect to the pairing (2.1). We will denote by $\mathcal{P}_{\pm}:V\to V_{\pm}$ the projections onto these respective subspaces.
    Let $\mathbb{G}_{\pm}$ denote the Lie subgroups of $\mathbb{G}$ associated with $\mathbb{g}_{\pm}$ so that (at least locally) $\mathbb{G}_{R}\simeq\mathbb{G}_{+}\times\mathbb{G}_{-}$. It will be useful to note for later that the coadjoint representation of $\mathbb{G}_{R}$ on $V\equiv\mathbb{g}^{\ast}$ can be expressed in terms of that of the subgroups $\mathbb{G}_{\pm}\subset\mathbb{G}$ as
    $\operatorname{Ad}^{R\ast}_{\varphi}\xi=\mathcal{P}_{-}(\operatorname{Ad}^{\ast}_{\varphi_{+}}\xi)+\mathcal{P}_{+}(\operatorname{Ad}^{\ast}_{\varphi_{-}}\xi)\,,$ (2.5)
@@
    A collection of independent $\operatorname{Ad}^{\ast}$-invariant functions on $\mathbb{g}^{\ast}$, say $H_{k}$, for $k=1,\dots,N$. They are in involution with respect to the Lie-Poisson bracket on $\mathbb{g}^{\ast}$ given by
    $\{f,g\}_{R}(\xi)=\big{\langle}\xi,[\nabla f(\xi)\,,\,\nabla g(\xi)]_{R}\big{\rangle}\,,\quad f,g\in C^{\infty}(\mathbb{g}^{\ast})$ (2.6)
    and generate a hierarchy of compatible flows on the Poisson manifold $\big{(}\mathbb{g}^{\ast},\{\,,\,\}_{R}\big{)}$ which take the form of Lax equations
@@
    When $\mathbb{g}$ and $V$ are subspaces of a common ambient Lie algebra and the coadjoint action $\operatorname{ad}^{\ast}:\mathbb{g}\times V\to V$ is given by a Lie bracket in this ambient Lie algebra, as will be the case for us in Section 3 below, the Lax equations (2.7) take the more familiar form
    $\partial_{t_{k}}L=[M_{k}(L),L]\,.$ (2.8)
    This is the content of Semenov-Tian-Shansky’s theorem, see, for instance, [STS].
The symplectic leaves of the Poisson manifold $\big{(}\mathbb{g}^{\ast},\{\,,\,\}_{R}\big{)}$ are given by the coadjoint orbits
${\cal O}_{\Lambda}=\{\operatorname{Ad}^{R\ast}_{\varphi}\Lambda;\varphi\in\mathbb{G}_{R}\}\,,\quad\text{for}\,\,\Lambda\in\mathbb{g}^{\ast}\,$ (2.9)
which therefore provide natural phase spaces for integrable systems. The key idea of [CDS] was to define a Lagrangian multiform on such a coadjoint orbit which produces the hierarchy (2.7) as its Euler-Lagrange equations. Specifically, let $L=\operatorname{Ad}^{R\ast}_{\varphi}\Lambda$ and define
$\mathscr{L}[\varphi]=\sum_{k=1}^{N}\,\mathscr{L}_{k}\,\mathrm{d}t_{k}\,,\quad\mathscr{L}_{k}=\big{\langle}\,L\,,\,\partial_{t_{k}}\varphi\cdot_{R}\varphi^{-1}\,\big{\rangle}-H_{k}(L)\,.$ (2.10)
Then, we have
Theorem 2.1 ([CDS]).
The Lagrangian $1$-form (2.10) satisfies the corner equations (1.4)-(1.5) of the multi-time Euler-Lagrange equations. The standard Euler-Lagrange equations (1.3) associated with the Lagrangian coefficients $\mathscr{L}_{k}$ give the hierarchy of compatible Lax equations (2.7)
$\partial_{t_{k}}L=\operatorname{ad}^{\ast}_{M_{k}(L)}L\,,\quad k=1,\dots,N\,.$ (2.11)
The closure relation [^2] The closure relation for the Lagrangian one-form $\mathscr{L}$ is, in fact, equivalent to the involutivity of the Hamiltonians $H_{k}$ with respect to the Lie-Poisson $R$-bracket {,}$\{\,\,,\,\}_{R}$. We do not discuss this result here and refer the interested reader to [CDS, Corollary 3.4]. holds: on solutions of (2.11) we have
$\partial_{t_{k}}\mathscr{L}_{j}-\partial_{t_{j}}\mathscr{L}_{k}=0\,,\quad j,k=1,\dots,N\,.$
In short, the machinery of [CDS] provides a systematic way of constructing a Lagrangian multiform for a hierarchy of Lax equations (2.7) on a coadjoint orbit (2.9). To apply this result to the cyclotomic Gaudin model we must identify its Lax matrix (1.17) as an element of an appropriate coadjoint orbit ${\cal O}_{\Lambda}\subset\mathbb{g}^{\ast}$, as in (2.9), for a suitable Lie algebra $\mathbb{g}$ and $R$-matrix. In particular, we will need to build a suitable model $V$ for the dual space $\mathbb{g}^{\ast}$ via a corresponding pairing $\langle\cdot,\cdot\rangle:V\times\mathbb{g}\to\mathbb{C}$. This will be the central goal of the next section.
@@ 3 Cyclotomic Gaudin model
The purpose of this section is to formulate a Lagrangian multiform for the cyclotomic Gaudin model using the machinery developped in [CDS]. As outlined at the end of Section 2, the first immediate task is to introduce the relevant infinite-dimensional Lie algebra $\mathbb{g}$ and a suitable model $V$ for its dual space $\mathbb{g}^{\ast}$ by constructing a nondegenerate invariant bilinear pairing as in (2.1). This will be achieved in Section 3.1, which culminates in a description of the relevant coadjoint orbits $\mathcal{O}_{\Lambda}\subset\mathbb{g}^{\ast}$ in Lemma 3.5. In Section 3.2 we specialise to the case of simple poles at $\omega^{k}\zeta_{r}$, $k\in\{0,\ldots,T-1\}$, $r\in\{1,\ldots,N\}$, and double poles at the origin and at infinity, and then describe the Lax matrix (1.17) of the cyclotomic Gaudin model as an element of this coadjoint orbit. In Section 3.3 we then describe the Lax equations as flows on this coadjoint orbit, as in (2.7). Finally, in Section 3.4 we put everything together to obtain the Lagrangian multiform for the cyclotomic Gaudin model.
@@ 3.1 Lie dialgebra framework
Let us fix $T\in\mathbb{Z}_{\geq 1}$ and pick a primitive $T$th root of unity $\omega\in\mathbb{C}^{\times}$. We can then define
$\Gamma\coloneqq\{1,\omega,\omega^{2},\ldots,\omega^{T-1}\}\,,$ (3.1)
a copy of the cyclic group $\mathbb{Z}_{T}$ of order $T$ that acts on $\mathbb{C}$ by multiplication. Further, we introduce the finite set of points $D=\{0,\zeta_{1},\ldots,\zeta_{N},\infty\}\subset\mathbb{C}P^{1}$ including the point at infinity such that the $\Gamma$-orbits of the points $\zeta_{1},\ldots,\zeta_{N}$ are pairwise disjoint, that is,
$\Gamma\zeta_{r}\cap\Gamma\zeta_{s}=\emptyset,\quad\text{for all}\,\,1\leq r\neq s\leq N\,.$ (3.2)
Note that unlike the non-zero finite points $\zeta_{1},\ldots,\zeta_{N}$, the origin and the point at infinity are fixed under the action of $\Gamma$.
Let $\mathfrak{g}$ be a finite-dimensional Lie algebra over $\mathbb{C}$ with an automorphism $\sigma$ of order $T$. For simplicity, we will only consider matrix Lie algebras in this work. The eigenspaces of $\sigma$,
$\mathfrak{g}^{(k)}=\{X\in\mathfrak{g}:\sigma(X)=\omega^{k}X\},\quad k\in\{0,\ldots,T-1\}\,,$ (3.3)
@@
$\mathfrak{g}=\bigoplus_{k=0}^{T-1}\mathfrak{g}^{(k)},\quad\text{with}\quad[\mathfrak{g}^{(k_{1})},\mathfrak{g}^{(k_{2})}]=\mathfrak{g}^{(k_{1}+k_{2}\,\text{mod}\,T)}\,.$ (3.4)
@@
$\lambda_{0}=\lambda\,,\qquad\lambda_{r}=\lambda-\zeta_{r}\quad\text{for}\,\,r\in\{1,\ldots,N\}\,,\qquad\lambda_{\infty}=\dfrac{1}{\lambda}\,.$ (3.5)
It will be convenient to also introduce an additional set of local parameters $\lambda_{r,k}=\lambda-\omega^{k}\zeta_{r}$ for all $k\in\{0,\ldots,T-1\}$ and $r\in\{1,\ldots,N\}$. We also introduce an index set $S=\{0,1,\ldots,N,\infty\}$ where 00 and $\infty$ denote labels for indices rather than points in $\mathbb{C}P^{1}$.
Let us denote by $\mathcal{F}_{D^{\prime}}$ the algebra of rational functions in the formal variable $\lambda$ with values in $\mathfrak{g}$ that are regular outside $D^{\prime}=\big{\{}\omega^{k}\zeta_{r}:k\in\{0,\ldots,T-1\},r\in\{1,\ldots,N\}\big{\}}\cup\{0,\infty\}$. Two subspaces of this algebra will be of relevance here: the subspace $\mathcal{F}_{D}^{\Gamma}$ of equivariant functions and the subspace $\Omega_{D}^{\Gamma}$ of equivariant one-forms that we define as
$\begin{split}&\mathcal{F}_{D}^{\Gamma}\coloneqq\{f\in\mathcal{F}_{D}(\mathfrak{g}):\sigma(f(\lambda))=f(\omega\lambda)\}\,,\\ &\Omega_{D}^{\Gamma}\coloneqq\{g\in\mathcal{F}_{D}(\mathfrak{g}):\sigma(g(\lambda))=\omega g(\omega\lambda)\}\,.\end{split}$ (3.6)
@@
$\begin{split}&f=\sum_{n=0}^{M_{0}-1}\frac{X_{0}^{(n)}}{\lambda^{n+1}}+\sum_{r=1}^{N}\sum_{k=0}^{T-1}\sum_{n=0}^{M_{r}-1}\frac{\sigma^{k}X_{r}^{(n)}}{(\omega^{-k}\lambda-\zeta_{r})^{n+1}}+\sum_{n=0}^{M_{\infty}}\frac{X_{\infty}^{(n)}}{\lambda_{\infty}^{n}}\,,\\ &g=\sum_{n=0}^{N_{0}-1}\frac{Y_{0}^{(n)}}{\lambda^{n+1}}+\sum_{r=1}^{N}\sum_{k=0}^{T-1}\sum_{n=0}^{N_{r}-1}\frac{\omega^{-k}\sigma^{k}Y_{r}^{(n)}}{(\omega^{-k}\lambda-\zeta_{r})^{n+1}}+\sum_{n=0}^{N_{\infty}}\frac{Y_{\infty}^{(n)}}{\lambda_{\infty}^{n}}\,,\end{split}$ (3.7)
@@
$\begin{split}&X_{0}^{(n)}\in\mathfrak{g}^{(-n-1)},\qquad X_{r}^{(n)}\in\mathfrak{g}\quad\text{for}\,\,r\in\{1,\ldots,N\},\qquad X_{\infty}^{(n)}\in\mathfrak{g}^{(n)}\,,\\ &Y_{0}^{(n)}\in\mathfrak{g}^{(-n)},\qquad Y_{r}^{(n)}\in\mathfrak{g}\quad\text{for}\,\,r\in\{1,\ldots,N\},\qquad Y_{\infty}^{(n)}\in\mathfrak{g}^{(n+1)}\,.\end{split}$ (3.8)
One can check that the subspace $\mathcal{F}_{D}^{\Gamma}$ is, in fact, a Lie subalgebra of $\mathcal{F}_{D^{\prime}}$. The subspace $\Omega_{D}^{\Gamma}$, on the other hand, contains the Lax matrix (1.17) of the model. To construct the cyclotomic Gaudin multiform à la [CDS], we will realise this Lax matrix as an element of a certain coadjoint orbit. To do so, let us define the loop algebra of formal Laurent series in the local parameter $\lambda_{r}$ with coefficients in $\mathfrak{g}$, for each $r\in S$,
@@
$[X\lambda_{r}^{i},Y\lambda_{r}^{j}]=[X,Y]\lambda_{r}^{i+j},\quad X,Y\in\mathfrak{g}\,,$ (3.10)
@@
$\widetilde{\mathfrak{g}}_{D}=\bigoplus_{r\in S}\widetilde{\mathfrak{g}}_{r}\,.$ (3.11)
@@
$[X,Y]_{r}(\lambda_{r})=[X_{r}(\lambda_{r}),Y_{r}(\lambda_{r})]\,.$ (3.12)

### v2

@@ 2 Lagrangian multiforms on coadjoint orbits
In this section, we briefly review the results of [11] where a Lagrangian multiform was obtained which produces a hierarchy of equations of motion in Lax form as its multi-time Euler–Lagrange equations. We will, in fact, be interested in applying this framework in the infinite-dimensional setting for which we refer more precisely to [11, Section 7]. In order to avoid confusion with the finite-dimensional Lie algebra $\mathfrak{g}$ used throughout the rest of this paper, in this section we denote by $\mathbb{g}$ the infinite-dimensional Lie algebra of interest. The construction of the specific Lie algebra $\mathbb{g}$ relevant for the cyclotomic Gaudin model will be the subject of the next section.
The framework of Lie dialgebras for constructing hierarchies of compatible Lax equations involves a number of ingredients. For the purpose of this paper, we will actually only require a particular instance of this general framework corresponding to the so-called Adler–Kostant–Symes scheme [2, 21, 38], so we will mainly focus on this case to ease the comparison with later sections. The ingredients are:
@@
    $\langle\cdot,\cdot\rangle\colon\ V\times\mathbb{g}\longrightarrow\mathbb{C}$ (2.1)
    with some representation $V$ of $\mathbb{G}$, allowing one to use $V$ as a model for the dual space $\mathbb{g}^{\ast}$ of $\mathbb{g}$. We will denote the representation of $\mathbb{G}$ by $\operatorname{Ad}^{\ast}\colon\mathbb{G}\times V\to V$, $(\varphi,\xi)\mapsto\operatorname{Ad}^{\ast}_{\varphi}\xi$ and the corresponding Lie algebra representation by $\operatorname{ad}^{\ast}\colon\mathbb{g}\times V\to V$, $(X,\xi)\mapsto\operatorname{ad}^{\ast}_{X}\xi$, and refer to these as the coadjoint representations.
    Note that if $\mathbb{g}$ were finite-dimensional then a good choice of representation space $V$ would be the algebraic dual $\mathbb{g}^{\prime}$ of $\mathbb{g}$, in which case (2.1) is given by the canonical pairing. In the infinite-dimensional setting we are considering, the algebraic dual is too big. When $\mathbb{g}$ is equipped with a nondegenerate invariant symmetric bilinear form $(\cdot,\cdot)\colon\mathbb{g}\times\mathbb{g}\to\mathbb{C}$ then one possible replacement for the algebraic dual is afforded by the smooth dual $(\mathbb{g},\cdot)\subset\mathbb{g}^{\prime}$ which is canonically isomorphic to $\mathbb{g}$ itself. However, in our present setting it turns out that the smooth dual will not be the appropriate notion of dual space, which is why we need the above more general working definition for the dual space $\mathbb{g}^{\ast}$.
@@
    A linear map $R\colon\mathbb{g}\to\mathbb{g}$ satisfying the modified classical Yang–Baxter equation (mCYBE)
    $[R(X),R(Y)]-R\left([R(X),Y]+[X,R(Y)]\right)=-[X,Y],\qquad\forall X,Y\in\mathbb{g}.$ (2.2)
[…]
@@
    into complementary subalgebras $\mathbb{g}_{\pm}\subset\mathbb{g}$. A solution to (2.2) is then given by $R=P_{+}-P_{-}$, where $P_{\pm}\colon\mathbb{g}\to\mathbb{g}_{\pm}$ denote the projections onto these respective subalgebras. In this case, the Lie algebra $\mathbb{g}_{R}$ is the direct sum of Lie algebras $\mathbb{g}_{R}=\mathbb{g}_{+}\oplus\mathbb{g}_{-}$. Now (2.3) induces a corresponding direct sum decomposition of the model $V$ for the dual space $\mathbb{g}^{\ast}$, namely
    $V=V_{+}\dotplus V_{-},$ (2.4)
    where $V_{\pm}$ is the orthogonal complement of $\mathbb{g}_{\pm}$ with respect to the pairing (2.1). We will denote by $\mathcal{P}_{\pm}\colon V\to V_{\pm}$ the projections onto these respective subspaces.
    Let $\mathbb{G}_{\pm}$ denote the Lie subgroups of $\mathbb{G}$ associated with $\mathbb{g}_{\pm}$, and $\mathbb{G}_{R}$ denote the Lie group associated with $\mathbb{g}_{R}$, so that (at least locally) $\mathbb{G}_{R}\simeq\mathbb{G}_{+}\times\mathbb{G}_{-}$. It will be useful to note for later that the coadjoint representation of $\mathbb{G}_{R}$ on $V\equiv\mathbb{g}^{\ast}$ can be expressed in terms of that of the subgroups $\mathbb{G}_{\pm}\subset\mathbb{G}$ as
    $\operatorname{Ad}^{R\ast}_{\varphi}\xi=\mathcal{P}_{-}\bigl{(}\operatorname{Ad}^{\ast}_{\varphi_{+}}\xi\bigr{)}+\mathcal{P}_{+}\bigl{(}\operatorname{Ad}^{\ast}_{\varphi_{-}}\xi\bigr{)}$ (2.5)
@@
    A collection of independent $\operatorname{Ad}^{\ast}$-invariant functions on $\mathbb{g}^{\ast}$, say $H_{k}$, for $k=1,\dots,N$. It follows from their $\operatorname{Ad}^{\ast}$-invariance that these functions are in involution with respect to the Lie–Poisson bracket on $\mathbb{g}^{\ast}$ given by
    $\{f,g\}_{R}(\xi)=\langle\xi,[\nabla f(\xi),\nabla g(\xi)]_{R}\rangle,\qquad f,g\in C^{\infty}(\mathbb{g}^{\ast}),$
    and generate a hierarchy of compatible flows on the Poisson manifold $(\mathbb{g}^{\ast},\{\,,\,\}_{R})$ which take the form of Lax equations
@@
    When $\mathbb{g}$ and $V$ are subspaces of a common ambient Lie algebra and the coadjoint action $\operatorname{ad}^{\ast}\colon\mathbb{g}\times V\to V$ is given by a Lie bracket in this ambient Lie algebra, as will be the case for us in Section 3 below, the Lax equations (2.6) take the more familiar form
    $\partial_{t_{k}}L=[M_{k}(L),L].$
    This is the content of Semenov-Tian-Shansky’s theorem, see, for instance, [30].
The symplectic leaves of the Poisson manifold $\bigl{(}\mathbb{g}^{\ast},\{\,,\,\}_{R}\bigr{)}$ are given by the coadjoint orbits
${\cal O}_{\Lambda}=\bigl{\{}\operatorname{Ad}^{R\ast}_{\varphi}\Lambda\mid\varphi\in\mathbb{G}_{R}\bigr{\}}\qquad\text{for}\quad\Lambda\in\mathbb{g}^{\ast},$ (2.7)
which therefore provide natural phase spaces for integrable systems. The key idea of [11] was to define a Lagrangian multiform on such a coadjoint orbit which produces the hierarchy (2.6) as its Euler–Lagrange equations. Specifically, let $L=\operatorname{Ad}^{R\ast}_{\varphi}\Lambda$ and define
$\mathscr{L}[\varphi]=\sum_{k=1}^{N}\mathscr{L}_{k}\,\mathrm{d}t_{k},\qquad\mathscr{L}_{k}=\bigl{\langle}L,\,\partial_{t_{k}}\varphi\cdot_{R}\varphi^{-1}\bigr{\rangle}-H_{k}(L),$ (2.8)
where $\cdot_{R}$ denotes the multiplication in $\mathbb{G}_{R}$. Then, we have the following.
Theorem 2.1 ([11]).
The Lagrangian $1$-form (2.8) satisfies the corner equations (1.2)–(1.3) of the multi-time Euler–Lagrange equations. The standard Euler–Lagrange equations (1.1) associated with the Lagrangian coefficients $\mathscr{L}_{k}$ give the hierarchy of compatible Lax equations (2.6)
$\partial_{t_{k}}L=\operatorname{ad}^{\ast}_{M_{k}(L)}L,\qquad k=1,\dots,N.$ (2.9)
The closure relation [^3] The closure relation for the Lagrangian one-form $\mathscr{L}$ is, in fact, equivalent to the involutivity of the Hamiltonians $H_{k}$ with respect to the Lie–Poisson $R$-bracket {,}$\{\,,\,\}_{R}$. We do not discuss this result here and refer the interested reader to [11, Corollary 3.4]. holds: on solutions of (2.9) we have
$\partial_{t_{k}}\mathscr{L}_{j}-\partial_{t_{j}}\mathscr{L}_{k}=0,\qquad j,k=1,\dots,N.$
In short, the machinery of [11] provides a systematic way of constructing a Lagrangian multiform for a hierarchy of Lax equations (2.6) on a coadjoint orbit (2.7). To apply this result to the cyclotomic Gaudin model we must identify its Lax matrix (1.10) as an element of an appropriate coadjoint orbit ${\cal O}_{\Lambda}\subset\mathbb{g}^{\ast}$, as in (2.7), for a suitable Lie algebra $\mathbb{g}$ and $R$-matrix. In particular, we will need to build a suitable model $V$ for the dual space $\mathbb{g}^{\ast}$ via a corresponding pairing $\langle\cdot,\cdot\rangle\colon V\times\mathbb{g}\to\mathbb{C}$. This will be the central goal of the next section.
@@ 3 Cyclotomic Gaudin model
The purpose of this section is to formulate a Lagrangian multiform for the cyclotomic Gaudin model using the machinery developed in [11]. As outlined at the end of Section 2, the first immediate task is to introduce the relevant infinite-dimensional Lie algebra $\mathbb{g}$ and a suitable model $V$ for its dual space $\mathbb{g}^{\ast}$ by constructing a nondegenerate invariant bilinear pairing as in (2.1). This will be achieved in Section 3.1, which culminates in a description of the relevant coadjoint orbits $\mathcal{O}_{\Lambda}\subset\mathbb{g}^{\ast}$ in Lemma 3.5. In Section 3.2, we specialise to the case of simple poles at $\omega^{k}\zeta_{r}$, $k\in\{0,\dots,T-1\}$, $r\in\{1,\dots,N\}$, and double poles at the origin and at infinity, and then describe the Lax matrix (1.10) of the cyclotomic Gaudin model as an element of this coadjoint orbit. In Section 3.3, we then describe the Lax equations as flows on this coadjoint orbit, as in (2.6). Finally, in Section 3.4, we put everything together to obtain the Lagrangian multiform for the cyclotomic Gaudin model.
@@ 3.1 Lie dialgebra framework
Let us fix $T\in\mathbb{Z}_{\geq 1}$ and pick a primitive $T$-th root of unity $\omega\in\mathbb{C}^{\times}$. We can then define
$\Gamma\coloneqq\bigl{\{}1,\omega,\omega^{2},\dots,\omega^{T-1}\bigr{\}},$
a copy of the cyclic group $\mathbb{Z}_{T}$ of order $T$ that acts on $\mathbb{C}$ by multiplication. Further, we introduce the finite set of points $D=\{0,\zeta_{1},\dots,\zeta_{N},\infty\}\subset\mathbb{C}P^{1}$ including the point at infinity such that the $\Gamma$-orbits of the points $\zeta_{1},\dots,\zeta_{N}$ are pairwise disjoint, that is,
$\Gamma\zeta_{r}\cap\Gamma\zeta_{s}=\varnothing\qquad\text{for all}\quad 1\leq r\neq s\leq N.$
Note that unlike the non-zero finite points $\zeta_{1},\dots,\zeta_{N}$, the origin and the point at infinity are fixed under the action of $\Gamma$.
Let $\mathfrak{g}$ be a finite-dimensional Lie algebra over $\mathbb{C}$ with an automorphism $\sigma$ of order $T$. For simplicity, we will only consider matrix Lie algebras in this work, with the Lie bracket being the commutator, and a nondegenerate invariant bilinear pairing given by the trace. We also assume that the automorphism $\sigma$ preserves the trace, that is, $\operatorname{Tr}(\sigma(x)\sigma(y))=\operatorname{Tr}(xy)$ for any $x,y\in\mathfrak{g}$. The eigenspaces of $\sigma$,
$\mathfrak{g}^{(k)}=\bigl{\{}X\in\mathfrak{g}\mid\sigma(X)=\omega^{k}X\bigr{\}},\qquad k\in\{0,\dots,T-1\},$ (3.1)
@@
$\mathfrak{g}=\mathfrak{g}^{(0)}\dotplus\dots\dotplus\mathfrak{g}^{(T-1)}\qquad\text{with}\quad\bigl{[}\mathfrak{g}^{(k_{1})},\mathfrak{g}^{(k_{2})}\bigr{]}=\mathfrak{g}^{(k_{1}+k_{2}\,\text{mod}\,T)}.$ (3.2)
@@
$\lambda_{0}=\lambda,\qquad\lambda_{r}=\lambda-\zeta_{r}\quad\text{for}\ r\in\{1,\dots,N\},\qquad\lambda_{\infty}=\dfrac{1}{\lambda}.$ (3.3)
It will be convenient to also introduce an additional set of local parameters $\lambda_{r,k}=\lambda-\omega^{k}\zeta_{r}$ for all $k\in\{0,\dots,T-1\}$ and $r\in\{1,\dots,N\}$. We also introduce an index set $S=\{0,1,\dots,N,\infty\}$ where 00 and $\infty$ denote labels for indices rather than points in $\mathbb{C}P^{1}$.
Let us denote by $\mathcal{F}_{D^{\prime}}$ the algebra of rational functions in the formal variable $\lambda$ with values in $\mathfrak{g}$ that are regular outside $D^{\prime}=\bigl{\{}\omega^{k}\zeta_{r}\mid k\in\{0,\dots,T-1\},\,r\in\{1,\dots,N\}\bigr{\}}\cup\{0,\infty\}$. Two subspaces of this algebra will be of relevance here: the subspace $\mathcal{F}_{D^{\prime}}^{\Gamma}$ of equivariant functions and the subspace $\Omega_{D^{\prime}}^{\Gamma}$ of equivariant one-forms that we define as
$\mathcal{F}_{D^{\prime}}^{\Gamma}\coloneqq\{f\in\mathcal{F}_{D^{\prime}}\mid\sigma(f(\lambda))=f(\omega\lambda)\},\qquad\Omega_{D^{\prime}}^{\Gamma}\coloneqq\{g\in\mathcal{F}_{D^{\prime}}\mid\sigma(g(\lambda))=\omega g(\omega\lambda)\}.$
@@
$f=\sum_{n=0}^{M_{0}-1}\frac{X_{0}^{(n)}}{\lambda^{n+1}}+\sum_{r=1}^{N}\sum_{k=0}^{T-1}\sum_{n=0}^{M_{r}-1}\frac{\sigma^{k}X_{r}^{(n)}}{\bigl{(}\omega^{-k}\lambda-\zeta_{r}\bigr{)}^{n+1}}+\sum_{n=0}^{M_{\infty}}\frac{X_{\infty}^{(n)}}{\lambda_{\infty}^{n}},$
$g=\sum_{n=0}^{N_{0}-1}\frac{Y_{0}^{(n)}}{\lambda^{n+1}}+\sum_{r=1}^{N}\sum_{k=0}^{T-1}\sum_{n=0}^{N_{r}-1}\frac{\omega^{-k}\sigma^{k}Y_{r}^{(n)}}{\bigl{(}\omega^{-k}\lambda-\zeta_{r}\bigr{)}^{n+1}}+\sum_{n=0}^{N_{\infty}}\frac{Y_{\infty}^{(n)}}{\lambda_{\infty}^{n}},$
@@
$X_{0}^{(n)}\in\mathfrak{g}^{(-n-1)},\qquad$ $X_{r}^{(n)}\in\mathfrak{g}\quad\text{for}\ r\in\{1,\dots,N\},\qquad$ $X_{\infty}^{(n)}\in\mathfrak{g}^{(n)},$
$Y_{0}^{(n)}\in\mathfrak{g}^{(-n)},\qquad$ $Y_{r}^{(n)}\in\mathfrak{g}\quad\text{for}\ r\in\{1,\dots,N\},\qquad$ $Y_{\infty}^{(n)}\in\mathfrak{g}^{(n+1)}.$ (3.4)
One can check that the subspace $\mathcal{F}_{D^{\prime}}^{\Gamma}$ is, in fact, a Lie subalgebra of $\mathcal{F}_{D^{\prime}}$. The subspace $\Omega_{D^{\prime}}^{\Gamma}$, on the other hand, contains the Lax matrix (1.10) of the model. To construct the cyclotomic Gaudin multiform à la [11], we will realise this Lax matrix as an element of a certain coadjoint orbit. To do so, let us define the loop algebra of formal Laurent series in the local parameter $\lambda_{r}$ with coefficients in $\mathfrak{g}$, for each $r\in S$,
@@
$\bigl{[}X\lambda_{r}^{i},Y\lambda_{r}^{j}\bigr{]}=[X,Y]\lambda_{r}^{i+j},\qquad X,Y\in\mathfrak{g},$ (3.6)
@@
$\widetilde{\mathfrak{g}}_{D}=\bigoplus_{r\in S}\widetilde{\mathfrak{g}}_{r}.$
@@
$[X,Y]_{r}(\lambda_{r})=[X_{r}(\lambda_{r}),Y_{r}(\lambda_{r})].$